# Enabling stratified sampling in high dimensions via nonlinear dimensionality reduction

### v1

@@ Abstract
We consider the problem of propagating the uncertainty from a possibly large number of random inputs through a computationally expensive model. Stratified sampling is a well-known variance reduction strategy, but its application, thus far, has focused on models with a limited number of inputs due to the challenges of creating uniform partitions in high dimensions. To overcome these challenges, we perform stratification with respect to the uniform distribution defined over the unit interval, and then derive the corresponding strata in the original space using nonlinear dimensionality reduction. We show that our approach is effective in high dimensions and can be used to further reduce the variance of multifidelity Monte Carlo estimators.
@@ 1 Introduction
Mathematical modeling and numerical simulations are fundamental in most scientific and engineering disciplines for advancing our ability to understand and predict complex phenomena [4]. Yet, the predictive power of these simulations is invariably affected by our imperfect knowledge of underlying mechanisms and their inherent variability. Ultimately, understanding and quantifying the uncertainty in computational model outputs is essential for establishing their validity. This led to increasing recent interest in the development of computational efficient strategies to propagate input uncertainty through complex computational models [27]. Many of these strategies consider uncertain inputs as random variables, and provide approximations for the expectation, or higher order moments, of one or multiple quantities of interests (QoIs). Consider the computational model $\mathcal{Q}\colon\mathbb{R}^{d}\to\mathbb{R}$, and let $X\sim\mu$ be a collection of input parameters with distribution $\mu$ on $\mathbb{R}^{d}$. We seek to estimate the quantity
@@
To make this computationally feasible, it is essential to reduce the variance of these estimates. Many methods have been proposed in the literature to achieve this. In this work, we focus on stratified sampling, which is based on a decomposition of the support of $X$ in multiple strata of smaller variance [1, Chapter 5]. This approach is known to scale poorly to high dimensions, as the number of partitions needed to keep a constant number of strata in each dimension grows exponentially. Variance reduction is also achieved by quasi-Monte Carlo estimators where random samples are replaced by deterministic low-discrepancy sequences [20], such as Halton, Hammersley, and Sobol’ sequences [10, 9, 26]. Even if quasi-Monte Carlo can theoretically achieve asymptotically faster convergence than standard Monte Carlo under appropriate regularity assumptions on the model, its performance may also suffer in high dimensions due to challenges of generating uncorrelated samples. Dimensionality reduction techniques have been applied to quasi-Monte Carlo methods to mitigate this problem, combined with smoothness in [21], and more recently with quadratic regression in [13]. We also mention importance sampling, where samples are drawn from a different distribution than the one of interest [15], and antithetic sampling where one aims to get samples that are negatively correlated [11]. Finally, an increasingly popular family of approaches include control variates [17], multilevel Monte Carlo [6], and multifidelity Monte Carlo, that leverage cheaper low-fidelity approximations of the expensive high-fidelity model [22]. For a comprehensive review on variance reduction techniques for Monte Carlo the interested reader is referred to [7, Chapter 4].
@@
We propose a methodology based on nonlinear dimensionality reduction to generate partitions that are adapted to the properties of the model $\mathcal{Q}$. In practice, we employ neural active manifolds (NeurAM) [28] to determine a one-dimensional manifold that follows the variability of the model, resulting in strata that tend to be separated by the level sets of $\mathcal{Q}$. NeurAM combines an autoencoder and a low-dimensional surrogate model built on a one-dimensional latent space. In addition, through projections on the inverse cumulative distribution function, it provides an invertible transformation between the latent space and a uniform distribution supported on the unit interval. Thus NeurAM allows a straightforward partition on the unit interval to be mapped to corresponding strata in the original domain. The fact that the stratification is performed in the one-dimensional unit interval is the crucial point allowing the method to scale to high-dimensional input domains. Since NeurAM and, more generally, dimensionality reduction, have already been successfully applied to improve the performance of multifidelity Monte Carlo estimators in [31, 30, 29, 19], we show how this novel stratification can also be implemented in the context of multifidelity estimators.
@@
Outline.
This paper proceeds as follows. In Section 2, we introduce our methodology, analyze the variance of the proposed estimator, and present a heuristic algorithm for stratification. Next, in Section 3 we apply this approach to multifidelity estimators. Then, in Section 4 we present numerical examples to demonstrate the properties and potential of our technique. Finally, Section 5 concludes the paper and suggests avenues for future research.
@@ 2 Stratified sampling
Under an appropriate allocation $\{N_{s}\}_{s=1}^{S}$ of the $N$ samples, which we will discuss in the next section, the variance of the stratified estimator is never larger than the variance of the corresponding standard Monte Carlo estimator with the same computational budget. However, the main challenge in stratified sampling is the selection of the strata $\{D_{s}\}_{s=1}^{S}$, which represents a serious obstacle for its application to high-dimensional models, since stratified sampling faces the curse of dimensionality [24, Lemma 3]. In high-dimensional settings, one can adopt a more effective approach known as Latin Hypercube Sampling (LHS), which involves drawing samples that are stratified across each dimension [18]. Intuitively, LHS is the high-dimensional analogue of placing one sample in each row and each column of a regular two-dimensional grid. Although LHS improves standard stratified sampling, its effectiveness also decreases for high-dimensional models. Moreover, this technique assumes that input variables are independent, making it unsuitable for problems with correlated inputs. A different strategy for constructing more clever strata has been proposed in [3], where the strata, defined as hyperrectangles, and their sample allocation are adaptively updated during the estimation process based on the directions that define these hyperrectangles. This approach enhances the accuracy of the estimators, particularly in the asymptotic regime where the number of samples and strata is large. However, since this method is adaptive, multiple iterations are required to achieve an effective stratification. Additionally, even after selecting the strata, it is not always straightforward to compute their probability, which is a necessary step for computing $\widehat{q}_{\mathrm{sMC}}$. To overcome these problems, we use nonlinear dimensionality reduction, and specifically, the NeurAM algorithm introduced in [28], to inform the stratification of the domain.
@@ 2.1 NeurAM-based stratification
$\mathcal{L}(\mathcal{E},\mathcal{D},\mathcal{S})$ $=\operatorname{\mathbb{E}}^{\mu}\left[(\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(\mathcal{D}(\mathcal{E}(X)))))^{2}\right]$ (2.5)
$\quad+\operatorname{\mathbb{E}}^{\mu}\left[(\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(X)))^{2}\right]$
@@
Note that a global optimum leading to a zero loss can be expressed in close form as
@@
where $\mathcal{I}$ stands for the identity function. This solution is however not easy to compute, due to the possibly large computational cost of evaluating the model $\mathcal{Q}=\mathcal{E}$, and the complexity of computing $\mathcal{D}$ as the right inverse of $\mathcal{Q}$. Therefore, in practice, we parameterize ~$\widetilde{\mathcal{E}}(\cdot,\mathfrak{e}),\widetilde{\mathcal{D}}(\cdot,\mathfrak{d}),\widetilde{\mathcal{S}}(\cdot,\mathfrak{s})$ as neural networks, and solve a minimization problem where the loss function is approximated by
~$\widetilde{\mathcal{L}}(\mathfrak{e},\mathfrak{d},\mathfrak{s})$ $=\frac{1}{M}\sum_{m=1}^{M}(\mathcal{Q}(x_{m})-\widetilde{\mathcal{S}}(\widetilde{\mathcal{E}}(\widetilde{\mathcal{D}}(\widetilde{\mathcal{E}}(x_{m};\mathfrak{e});\mathfrak{d});\mathfrak{e});\mathfrak{s}))^{2}$ (2.7)
$\quad+\frac{1}{M}\sum_{m=1}^{M}(\mathcal{Q}(x_{m})-\widetilde{\mathcal{S}}(\widetilde{\mathcal{E}}(x_{m};\mathfrak{e});\mathfrak{s}))^{2}$
@@
()#$(\mathcal{F}\circ\mathcal{E})_{\#}\mu=\mathcal{U}([0,1]),$ (2.9)
@@
Similarly to the autoencoder and the surrogate, also the CDF $\mathcal{F}$ must be approximated. We propose to use the empirical distribution $\widetilde{\mathcal{F}}$ as represented by the histogram of {~$\{\widetilde{\mathcal{E}}(x_{k};\mathfrak{e})\}_{k=1}^{K}$ where $x_{k}\sim\mu$. Notice that $K$ is not limited by the available computational budget, since the cost of evaluating the encoder is negligible. For additional details on the construction of the neural active manifold we refer to [28].
@@
Therefore, due to Lemma 2.1, the proposed NeurAM-based stratified estimator and its variance become
@@
$\operatorname{Var}[\widehat{q}_{\mathrm{sMC}}]$ $=\sum_{s=1}^{S}\frac{\lambda(A_{s})^{2}}{N_{s}}\operatorname{Var}^{\mu}[\mathcal{Q}(X)|X\in D_{s}].$
@@
The samples $x_{n}^{(s)}$ in the stratum $D_{s}$ can be obtained by sampling from the probability distribution $\mu$ and by rejecting the values for which $\mathcal{F}(\mathcal{E}(x_{n}^{(s)}))\not\in A_{s}$.
@@ 2.2 Allocation strategies and analysis of the variance
which shows that the variance reduction is greater when the differences between the global expectation and the local expectations within each stratum are larger.
Figure 1: Comparison between the NeurAM-based stratification (left) and the standard stratification made with a regular grid (right), for the simple linear model in Example 2.4.
@@
$\mathcal{F}(t)=\begin{cases}0,&\text{ if }t\leq-2,\\ \frac{1}{8}t^{2}+\frac{1}{2}t+\frac{1}{2},&\text{ if }-2\leq t\leq 0,\\ -\frac{1}{8}t^{2}+\frac{1}{2}t+\frac{1}{2},&\text{ if }0\leq t\leq 2,\\ 1,&\text{ if }t\geq 2.\end{cases}$ (2.23)
@@
From equation (2.12), the NeurAM-based stratification of $[-1,1]^{2}$ is given by
$D_{1}$ =$=\left\{x\in[-1,1]^{2}\colon-1<x_{1}<\sqrt{2}-1,\quad-1<x_{2}<-x_{1}+\sqrt{2}-2\right\},$ (2.25)
$D_{2}$ =$=\left\{x\in[-1,1]^{2}\colon-1<x_{1}<1,\quad-x_{1}+\sqrt{2}-2<x_{2}<-x_{1}\right\},$
$D_{3}$ =$=\left\{x\in[-1,1]^{2}\colon-1<x_{1}<1,\quad-x_{1}<x_{2}<-x_{1}+2-\sqrt{2}\right\},$
$D_{4}$ =$=\left\{x\in[-1,1]^{2}\colon 1-\sqrt{2}<x_{1}<1,\quad-x_{1}+2-\sqrt{2}<x_{2}<1\right\},$
which is shown in the left plot in Figure 1. We then have $\lambda(A_{s})=1/4$ for all $s=1,\dots,4$, and the local variances (we omit their derivation for brevity) are
@@
which is shown in the right plot in Figure 1. In this case, for all $s=1,\dots,4$ we have

### v2

@@ Abstract
We consider the problem of propagating the uncertainty from a possibly large number of random inputs through a computationally expensive model. Stratified sampling is a well-known variance reduction strategy, but its application, thus far, has focused on models with a limited number of inputs due to the challenges of creating uniform partitions in high dimensions. To overcome these challenges, we propose a simple methodology for constructing an effective stratification of the input domain that is adapted to the model response. Our approach leverages neural active manifolds, a recently introduced nonlinear dimensionality reduction technique based on neural networks that identifies a one-dimensional manifold capturing most of the model variability. The resulting one-dimensional latent space is mapped to the unit interval, where stratification is performed with respect to the uniform distribution. The corresponding strata in the original input space are then recovered through the neural active manifold, generating partitions that tend to follow the level sets of the model. We show that our approach is effective in high dimensions and can be used to further reduce the variance of multifidelity Monte Carlo estimators.
@@ 1 Introduction
Mathematical modeling and numerical simulations are fundamental in most scientific and engineering disciplines for advancing our ability to understand and predict complex phenomena [FoM21]. Yet, the predictive power of these simulations is invariably affected by our imperfect knowledge of underlying mechanisms and their inherent variability. Ultimately, understanding and quantifying the uncertainty in computational model outputs is essential for establishing their validity. This led to increasing recent interest in the development of computational efficient strategies to propagate input uncertainty through complex computational models [Sul15]. Many of these strategies consider uncertain inputs as random variables, and provide approximations for the expectation, or higher order moments, of one or multiple quantities of interests (QoIs). Consider the computational model $\mathcal{Q}\colon\mathbb{R}^{d}\to\mathbb{R}$, and let $X\sim\mu$ be a collection of input parameters with distribution $\mu$ on $\mathbb{R}^{d}$. We seek to estimate the quantity
@@
To make this computationally feasible, it is essential to reduce the variance of these estimates. Many methods have been proposed in the literature to achieve this. In this work, we focus on stratified sampling, which is based on a decomposition of the support of $X$ in multiple strata of smaller variance [Coc77, Chapter 5]. This approach is known to scale poorly to high dimensions, as the number of partitions needed to keep a constant number of strata in each dimension grows exponentially. Variance reduction is also achieved by quasi-Monte Carlo estimators where random samples are replaced by deterministic low-discrepancy sequences [MoC95], such as Halton, Hammersley, and Sobol’ sequences [HaH65, Hal60, Sob67]. Even if quasi-Monte Carlo (qMC) can theoretically achieve asymptotically faster convergence than standard Monte Carlo under appropriate regularity assumptions on the model, its performance may also suffer in high dimensions due to challenges of generating uncorrelated samples. Furthermore, qMC requires the use of a power-of-two number of samples to maintain optimal low-discrepancy properties. Dimensionality reduction techniques have been applied to quasi-Monte Carlo methods to mitigate this problem, combined with smoothness in [MoC96], and more recently with quadratic regression in [ImT25]. We also mention importance sampling, where samples are drawn from a different distribution than the one of interest [KlV78], and antithetic sampling where one aims to get samples that are negatively correlated [HaM56]. Finally, an increasingly popular family of approaches include control variates [Lem17], multilevel Monte Carlo [Gil15], and multifidelity Monte Carlo, that leverage cheaper low-fidelity approximations of the expensive high-fidelity model [NgW14]. For a comprehensive review on variance reduction techniques for Monte Carlo the interested reader is referred to [Gla04, Chapter 4].
@@
We propose a methodology based on nonlinear dimensionality reduction to generate partitions that are adapted to the properties of the model $\mathcal{Q}$. In practice, we employ neural active manifolds (NeurAM) [ZGS25] to determine a one-dimensional manifold that follows the variability of the model, resulting in strata that tend to be separated by the level sets of $\mathcal{Q}$. NeurAM combines an autoencoder and a low-dimensional surrogate model built on a one-dimensional latent space. In addition, through projections on the inverse cumulative distribution function, it provides an invertible transformation between the latent space and a uniform distribution supported on the unit interval. Thus NeurAM allows a straightforward partition on the unit interval to be mapped to corresponding strata in the original domain. The fact that the stratification is performed in the one-dimensional unit interval is the crucial point allowing the method to scale to high-dimensional input domains. Since NeurAM and, more generally, dimensionality reduction, have already been successfully applied to improve the performance of multifidelity Monte Carlo estimators in [ZGE23, ZGS24a, ZGS24b, MZK24], we show how this novel stratification can also be implemented in the context of multifidelity estimators.
@@
Outline
This paper proceeds as follows. In Section˜2, we introduce our methodology, analyze the variance of the proposed estimator, and present a heuristic algorithm for stratification. Next, in Section˜3 we apply this approach to multifidelity estimators. Then, in Section˜4 we present numerical examples to demonstrate the properties and potential of our technique. Finally, Section˜5 concludes the paper and suggests avenues for future research.
@@ 2 Stratified sampling
Under an appropriate allocation $\{N_{s}\}_{s=1}^{S}$ of the $N$ samples, which we will discuss in the next section, the variance of the stratified estimator is never larger than the variance of the corresponding standard Monte Carlo estimator with the same computational budget. However, the main challenge in stratified sampling is the selection of the strata $\{D_{s}\}_{s=1}^{S}$, which represents a serious obstacle for its application to high-dimensional models, since stratified sampling faces the curse of dimensionality [PeK22, Lemma 3].
In high-dimensional settings, one can adopt a more effective approach known as Latin Hypercube Sampling (LHS), which involves drawing samples that are stratified across each dimension [MBC79]. Intuitively, LHS is the high-dimensional analogue of placing one sample in each row and each column of a regular two-dimensional grid. Although LHS improves standard stratified sampling, its effectiveness also decreases for high-dimensional models. Moreover, this technique assumes that input variables are independent, making it unsuitable for problems with correlated inputs. A different strategy for constructing more clever strata has been proposed in [EFJ11], where the strata, defined as hyperrectangles, and their sample allocation are adaptively updated during the estimation process based on the directions that define these hyperrectangles. This approach enhances the accuracy of the estimators, particularly in the asymptotic regime where the number of samples and strata is large. Alternative strategies are based on iterative refinements of an existing stratification. In [STH15], new samples are added sequentially by dividing existing strata along the direction identified by the unstratified component of largest magnitude. This approach is further combined with hierarchical LHS in [Shi16] to improve its performance for high-dimensional problems. In [PeK22], the authors propose to combine optimal and proportional sampling, in a way that is applicable to nonsmooth models, and provide theoretical estimates for the expected performance. LHS can also be applied within each stratum to further reduce the variance of the resulting stratified estimator [KrP24]. In all these adaptive strategies, multiple iterations are required to achieve an effective stratification, and new samples need to be generated at each iteration.
Additionally, even after selecting the strata, it is not always straightforward to compute their probability, which is a necessary step for computing $\widehat{q}_{\mathrm{sMC}}$. To overcome these problems, we use nonlinear dimensionality reduction, and specifically, the NeurAM algorithm introduced in [ZGS25], to inform the stratification of the domain.
@@ 2.1 NeurAM-based stratification
$\mathcal{L}(\mathcal{E},\mathcal{D},\mathcal{S})$ $=\operatorname{\mathbb{E}}^{\mu}\left[(\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(\mathcal{D}(\mathcal{E}(X)))))^{2}\right]+\operatorname{\mathbb{E}}^{\mu}\left[(\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(X)))^{2}\right]$ (2.5)
@@
The three terms in the loss function play different roles. Specifically, the second term guarantees that the surrogate model $\mathcal{S}$, defined on the latent space of the autoencoder, accurately approximates the original model $\mathcal{Q}$, i.e., $\mathcal{Q}\simeq\mathcal{S}\circ\mathcal{E}$. The first term, which is instead essential to identify the NeurAM, enforces that projecting a point onto such one-dimensional manifold should preserve the corresponding model output, i.e., $Q(X~)\mathcal$$\simeq Q(\widetilde{X})$, where $\widetilde{X}=\mathcal{D}(\mathcal{E}(X))$. Since evaluating $\mathcal{Q}$ directly can be expensive, the surrogate model is used to approximate $\mathcal{Q}(X)$ at the projection $\widetilde{X}$ along the NeurAM. Finally, the third term ensures that if a point already lies on the manifold, then it is projected onto itself, i.e., $\mathcal{D}(\mathcal{E}(\widetilde{X}))\simeq\widetilde{X}$, or equivalently, that the NeurAM is made up of fixed points of the composition of the encoder and decoder. Note that a global optimum leading to a zero loss can be expressed in closed form as
@@
where $\mathcal{I}$ stands for the identity function. This solution is however not easy to compute, due to the possibly large computational cost of evaluating the model $\mathcal{Q}=\mathcal{E}$, and the complexity of computing $\mathcal{D}$ as the right inverse of $\mathcal{Q}$. Therefore, in practice, we parameterize $\widetilde{\mathcal{E}}(\cdot,\mathfrak{e}),\widetilde{\mathcal{D}}(\cdot,\mathfrak{d}),\widetilde{\mathcal{S}}(\cdot,\mathfrak{s})$ as neural networks with weights $\mathfrak{e},\mathfrak{d},\mathfrak{s}$, respectively, and solve a minimization problem where the loss function is approximated by
$\widetilde{\mathcal{L}}(\mathfrak{e},\mathfrak{d},\mathfrak{s})$ $=\frac{1}{M}\sum_{m=1}^{M}(\mathcal{Q}(x_{m})-\widetilde{\mathcal{S}}(\widetilde{\mathcal{E}}(\widetilde{\mathcal{D}}(\widetilde{\mathcal{E}}(x_{m};\mathfrak{e});\mathfrak{d});\mathfrak{e});\mathfrak{s}))^{2}+\frac{1}{M}\sum_{m=1}^{M}(\mathcal{Q}(x_{m})-\widetilde{\mathcal{S}}(\widetilde{\mathcal{E}}(x_{m};\mathfrak{e});\mathfrak{s}))^{2}$ (2.7)
@@
$(\mathcal{F}\circ\mathcal{E})_{\#}\mu=\mathcal{U}([0,1]),$ (2.9)
@@
Similarly to the autoencoder and the surrogate, also the CDF $\mathcal{F}$ must be approximated. We propose to use the empirical distribution $\widetilde{\mathcal{F}}$ as represented by the histogram of $\{\widetilde{\mathcal{E}}(x_{k};\mathfrak{e})\}_{k=1}^{K}$ where $x_{k}\sim\mu$. Notice that $K$ is not limited by the available computational budget, since the cost of evaluating the encoder is negligible. For additional details on the construction of the neural active manifold we refer to [ZGS25].
@@
Therefore, due to Lemma˜2.1, the proposed NeurAM-based stratified estimator and its variance become
@@
$\operatorname{Var}[\widehat{q}_{\mathrm{sMC}}]$ $=\sum_{s=1}^{S}\frac{\lambda(A_{s})^{2}}{N_{s}}\operatorname{Var}^{\mu}[\mathcal{Q}(X)|X\in D_{s}],$
where the second expression enables direct computation of the estimator variance by approximating the variances within each stratum. The main steps to construct the stratified estimator are summarized in Algorithm˜1.
@@
The samples $x_{n}^{(s)}$ in the stratum $D_{s}$ can be obtained by sampling from the probability distribution $\mu$ and by rejecting the values for which $\mathcal{F}(\mathcal{E}(x_{n}^{(s)}))\not\in A_{s}$. By construction, this guarantees that the accepted sample $x_{n}^{(s)}$ always lies within the domain $\mathbb{D}$, since all candidates are sampled from $\mu$. Moreover, since acceptance is conditioned on $\mathcal{F}(\mathcal{E}(x_{n}^{(s)}))\in A_{s}$, the resulting sample necessarily belongs to $D_{s}$ by definition. Therefore, it is not possible to produce samples outside the support of $\mathbb{D}$, or failing to cover portions of $\mathbb{D}$. In practice, this also means that the method can be used with existing datasets without requiring a specialized sampling procedure.
@@
Input: Model $\mathcal{Q}$
Input distribution $\mu$ with support $\mathbb{D}$ or samples from it
Number of strata $S$ and stratification of the unit interval $\{a_{s}\}_{s=0}^{S}$
Computational budget $N$ and allocation strategy $\{N_{s}\}_{s=1}^{S}$
Output: Estimator $\widehat{q}_{\mathrm{sMC}}$ of $q=\operatorname{\mathbb{E}}^{\mu}[\mathcal{Q}(X)]$
  1. 1:
    Construct the NeurAM $(\mathcal{E},\mathcal{D},\mathcal{S})$ by minimizing the loss function $\mathcal{L}$ in equation (2.5)
  2. 2:
    Construct the CDF $\mathcal{F}$ from equation (2.8)
  3. 3:
    Define the subdomains $\{A_{s}\}_{s=1}^{S}$ as $A_{s}=[a_{s-1},a_{s}]$
  4. 4:
    Compute the coefficients $\{\lambda(A_{s})\}_{s=1}^{S}$ as $\lambda(A_{s})=a_{s}-a_{s-1}$
  5. 5:
    Define the subdomains $\{D_{s}\}_{s=1}^{S}$ using equation (2.12)
  6. 6:
    Draw the samples $\{\{x_{n}^{(s)}\}_{n=1}^{N_{s}}\}_{s=1}^{S}$ such that $x_{n}^{(s)}\sim\mu|_{D_{s}}$ as explained in Remark˜2.2
  7. 7:
    Compute the estimator $\widehat{q}_{\mathrm{sMC}}$ using equation (2.15)
Algorithm 1 NeurAM-based stratified Monte Carlo estimator
@@ 2.2 Allocation strategies and analysis of the variance
which shows that the variance reduction is greater when the differences between the global expectation and the local expectations within each stratum are larger. We also mention that hybrid allocation strategies that use both proportional and optimal allocation have been introduced in [PeK22].
Figure 1: Comparison between the NeurAM-based stratification (left) and the standard stratification made with a regular grid (right), for the simple linear model in Example˜2.4.
@@
[…]
@@
From equation (2.12), the NeurAM-based stratification of $1,1]^{2}$ is given by
$D_{1}$ $=\left\{x\in[-1,1]^{2}\colon-1<x_{1}<\sqrt{2}-1,\quad-1<x_{2}<-x_{1}+\sqrt{2}-2\right\},$ (2.25)
$D_{2}$ $=\left\{x\in[-1,1]^{2}\colon-1<x_{1}<1,\quad-x_{1}+\sqrt{2}-2<x_{2}<-x_{1}\right\},$
$D_{3}$ $=\left\{x\in[-1,1]^{2}\colon-1<x_{1}<1,\quad-x_{1}<x_{2}<-x_{1}+2-\sqrt{2}\right\},$
$D_{4}$ $=\left\{x\in[-1,1]^{2}\colon 1-\sqrt{2}<x_{1}<1,\quad-x_{1}+2-\sqrt{2}<x_{2}<1\right\},$
which is shown in the left plot in Figure˜1. We then have $\lambda(A_{s})=1/4$ for all $s=1,\dots,4$, and the local variances (we omit their derivation for brevity) are
@@
which is shown in the right plot in Figure˜1. In this case, for all $s=1,\dots,4$ we have
@@
In the following result, we quantify how the accuracy of NeurAM, i.e., the magnitude of the achieved minimum of the loss function $\mathcal{L}$ in equation (2.5), influences the variance of the stratified Monte Carlo estimator.
Theorem 2.5.
Let $\mathcal{E},\mathcal{D},\mathcal{S}$ satisfy $\mathcal{L}(\mathcal{E},\mathcal{D},\mathcal{S})=\varepsilon\geq 0$, where $\mathcal{L}$ is defined in equation (2.5), and assume $\mathcal{F}$ to be invertible. Then, it holds for all $s=1,\dots,S$
$\operatorname{Var}^{\mu}[\mathcal{Q}(X)|X\in D_{s}]\leq 2\operatorname{Var}^{\mathcal{U}}[\mathcal{S}(\mathcal{F}^{-1}(U))|U\in A_{s}]+\frac{2\varepsilon}{\lambda(A_{s})}.$ (2.31)
Moreover, we have
$\operatorname{Var}[\widehat{q}_{\mathrm{sMC}}]\leq 2\sum_{s=1}^{S}\frac{\lambda(A_{s})^{2}}{N_{s}}\operatorname{Var}^{\mu}[\mathcal{S}(\mathcal{F}^{-1}(U))|U\in A_{s}]+2\varepsilon\sum_{s=1}^{S}\frac{\lambda(A_{s})}{N_{s}}.$ (2.32)
Proof.
By Cauchy–Schwarz inequality and properties of conditional expectations, we have
$\operatorname{Var}^{\mu}[\mathcal{Q}(X)|X\in D_{s}]$ $\leq 2\left(\operatorname{Var}^{\mu}[\mathcal{S}(\mathcal{E}(X))|X\in D_{s}]+\operatorname{Var}^{\mu}[\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(X))|X\in D_{s}]\right)$ (2.33)
$\leq 2\left(\operatorname{Var}^{\mu}[\mathcal{S}(\mathcal{E}(X))|X\in D_{s}]+\operatorname{\mathbb{E}}^{\mu}[(\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(X)))^{2}|X\in D_{s}]\right)$
$\leq 2\left(\operatorname{Var}^{\mu}[\mathcal{S}(\mathcal{E}(X))|X\in D_{s}]+\frac{\operatorname{\mathbb{E}}^{\mu}[(\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(X)))^{2}]}{\mu(D_{s})}\right).$
Then, notice that by definition of $D_{s}$ and Lemma˜2.1, we get $\mu(D_{s})=\lambda(A_{s})$ and
$\operatorname{Var}^{\mu}[\mathcal{S}(\mathcal{E}(X))|X\in D_{s}]$ $=\operatorname{Var}^{\mu}[\mathcal{S}(\mathcal{F}^{-1}(\mathcal{F}(\mathcal{E}(X))))|\mathcal{F}(\mathcal{E}(X))\in A_{s}]$ (2.34)
$=\operatorname{Var}^{\mathcal{U}}[\mathcal{S}(\mathcal{F}^{-1}(U))|U\in A_{s}],$
which, together with the fact that
$\operatorname{\mathbb{E}}^{\mu}[(\mathcal{Q}(X)-\mathcal{S}(\mathcal{E}(X)))^{2}]\leq\mathcal{L}(\mathcal{E},\mathcal{D},\mathcal{S})=\varepsilon,$ (2.35)
implies the first desired result. The second estimate follows directly from the first by equation (2.15). ∎
Applying Theorem˜2.5, we can obtain upper bounds for the variances of the stratified Monte Carlo estimators $\widehat{q}_{\mathrm{sMC}}^{(1)}$ and $\widehat{q}_{\mathrm{sMC}}^{(2)}$ with optimal and proportional allocation strategies, respectively. Specifically, using equation (2.19), we deduce
$\operatorname{Var}[\widehat{q}_{\mathrm{sMC}}^{(1)}]$ $\leq\frac{2}{N}\left\{\sum_{s=1}^{S}\lambda(A_{s})\sqrt{\operatorname{Var}^{\mathcal{U}}[\mathcal{S}(\mathcal{F}^{-1}(U))|U\in A_{s}]}\right\}^{2}+\frac{2S\varepsilon}{N},$ (2.36)
$\operatorname{Var}[\widehat{q}_{\mathrm{sMC}}^{(2)}]$ $\leq\frac{2}{N}\sum_{s=1}^{S}\lambda(A_{s})\operatorname{Var}^{\mathcal{U}}[[\mathcal{S}(\mathcal{F}^{-1}(U))|U\in A_{s}]+\frac{2S\varepsilon}{N}.$
In Theorem˜2.5, and consequently in equation (2.36), we see that an additional term appears that depends on the accuracy of the NeurAM. On the other hand, the first term reduces, up to the constant factor $2$, to the variance of the stratified estimator for the one-dimensional model $\mathcal{S}\circ\mathcal{F}^{-1}$ with respect to the uniform distribution $\mathcal{U}([0,1])$. This allows us to apply existing theoretical results for one-dimensional stratification. In particular, we obtain the following result.
Corollary 2.6.
Let $\mathcal{E},\mathcal{D},\mathcal{S}$ satisfy $\mathcal{L}(\mathcal{E},\mathcal{D},\mathcal{S})=\varepsilon\geq 0$, where $\mathcal{L}$ is defined in equation (2.5), and assume $\mathcal{F}$ to be invertible and $\mathcal{S}\circ\mathcal{F}^{-1}$ to be Lipschitz continuous. Using the uniform stratification of Example˜2.3, it holds
$\operatorname{Var}[\widehat{q}_{\mathrm{sMC}}^{\mathrm{u},(i)}]\leq\frac{\left\|(\mathcal{S}\circ\mathcal{F}^{-1})^{\prime}\right\|^{2}_{L^{\infty}([0,1])}}{6S^{2}N}+\frac{2S\varepsilon}{N},$ (2.37)
for both $i=1,2$, i.e., optimal and proportional allocations, respectively.
Proof.
Following [AsG07, Chapter V, Example 7.1], we have
$\operatorname{Var}^{\mathcal{U}}[[\mathcal{S}(\mathcal{F}^{-1}(U))|U\in A_{s}]\leq\frac{\left\|(\mathcal{S}\circ\mathcal{F}^{-1})^{\prime}\right\|^{2}_{L^{\infty}([0,1])}}{12S^{2}},$ (2.38)
which, together with equation (2.36), yields the desired result. ∎
Remark 2.7.
Corollary˜2.6 shows that the variance scales as $\mathcal{O}(S^{-2})$, but only up to the point where the error introduced by the NeurAM, quantified by $\varepsilon$, becomes dominant. Beyond this point, further refining the stratification may not reduce the variance and can even cause it to increase. By Corollary˜2.6, the optimal number of strata is
$S^{*}=\min\left\{\left\lceil\left(\frac{\left\|(\mathcal{S}\circ\mathcal{F}^{-1})^{\prime}\right\|^{2}_{L^{\infty}([0,1])}}{6\varepsilon}\right)^{1/3}\right\rceil,\;N\right\},$ (2.39)
which highlights that increasing the stratification is beneficial only when the NeurAM approximation is sufficiently accurate. Moreover, if we want the variance to vanish as $N\to\infty$, then even when $\varepsilon\ll 1$, we cannot choose $S=N$, since the error term would remain of order $\mathcal{O}(\varepsilon)$. We finally remark that this analysis is independent of the dimension $d$ of the domain $\mathbb{D}$ of the model $\mathcal{Q}$, which is one of the main advantages of this approach.